OpenStudy (anonymous):

find the area of the region bounded by y=2x+2, y=2/3x-2, and the a-axis

OpenStudy (anonymous):

Is that second curve the line \(y=\dfrac{2}{3}x-2\) or the curve \(y=\dfrac{2}{3x-2}\)?